OpenStudy (mamiinikki):

27/35 divided by 9/4

Directrix (directrix):

Invert the divisor and multiply: 27/35 divided by 9/4 = 27 4 ---- * ------- = 35 9

Directrix (directrix):

Simplify the answer.

OpenStudy (skullpatrol):

$$\Huge\dfrac{a}{b}\div\dfrac{c}{d}=\dfrac{a}{b}\cdot\dfrac{d}{c}$$ $$\Huge\dfrac{27}{35}\div\dfrac{9}{4}=\dfrac{27}{35}\cdot\dfrac{4}{9}$$ $$\Huge=\dfrac{\cancel9\cdot 3}{35}\cdot\dfrac{4}{\cancel9}$$ can you finish the rest?